OpenStudy (anonymous):

Evaluate the function g(t) = 3/2t - 5/2, when t = 3. A. 6½ B. 14 C. 2 D. 6

OpenStudy (anonymous):

C

OpenStudy (anonymous):

is it \[g(t)=\frac{3}{2t}+\frac{5}{2}\]??

OpenStudy (anonymous):

\[g(t)=\frac{3t}{2}-\frac{5}{2}\]??